OpenStudy (anonymous):

Find f(2) and f(a + h) when f(x) = 3x^2 + 2x + 4 f(2)= 3(2)^2+2(2)+4 = 20 but i dont get the f(a+h) f(a+h)= 3(a+h)^2+2(a+h)+4 right?

OpenStudy (anonymous):

subtitution f=2 and f=(a+h)

OpenStudy (anonymous):

f(2)= 3(2)^2+2(2)+4 f(a+h)=3(a+h)^2+2(a+h)+4 you can solve from here

OpenStudy (anonymous):

@Nancy_Lam i already did f(2) i need help with simplifying the second one: 3(a+h)^2+2(a+h)+4

OpenStudy (anonymous):

3(a+h)^2+2(a+h)+4 is easy to solve

OpenStudy (anonymous):

3a+3h^2+2a+2h+4?

OpenStudy (anonymous):

[3*(a+h)(a+h)] + [2a+2h] + 4 [3*( a^2+2ah+h^2)] + 2a+ 2h +4 can you solve from here?

OpenStudy (anonymous):

is it 3a^2+2ah+2a+h^2+2h+4?

OpenStudy (anonymous):

= 3a^2+3h^2+6ah+2a+2h+4

OpenStudy (anonymous):

oh ok thank you! :)

OpenStudy (anonymous):

nedal?